# Generalization under Byzantine & Poisoning Attacks: Tight Stability Bounds in Robust Distributed Learning

### v1

@@ Abstract
Robust distributed learning algorithms aim to maintain good performance in distributed and federated settings, even in the presence of misbehaving workers. Two primary threat models have been studied: Byzantine attacks, where misbehaving workers can send arbitrarily corrupted updates, and data poisoning attacks, where misbehavior is limited to manipulation of local training data. While prior work has shown comparable optimization error under both threat models, a fundamental question remains open: How do these threat models impact generalization? Empirical evidence suggests a gap between the two threat models, yet it remains unclear whether it is fundamental or merely an artifact of suboptimal attacks. In this work, we present the first theoretical investigation into this problem, formally showing that Byzantine attacks are intrinsically more harmful to generalization than data poisoning. Specifically, we prove that: (i) under data poisoning, the uniform algorithmic stability of a robust distributed learning algorithm, with optimal optimization error, degrades by an additive factor of $\varTheta(\frac{f}{n-f})$, with $f$ the number of misbehaving workers out of $n$; and (ii) In contrast, under Byzantine attacks, the degradation is in $\mathcal{O}\big{(}\sqrt{\frac{f}{n-2f}}\big{)}$. This difference in stability leads to a generalization error gap that is especially significant as $f$ approaches its maximum value $\frac{n}{2}$.
@@ 1 Introduction
With the proliferation of large-scale data and pervasive connectivity, distributed learning has evolved from a theoretical concept into a fundamental paradigm in modern AI [24]. From federated smartphones to geo-distributed datacenters, algorithms now orchestrate millions of workers toward shared objectives. Yet this democratization of learning comes at a cost: unlike the idealized setting of perfectly cooperative workers, real-world systems face a range of adversarial behaviors, broadly categorized under the umbrella of Byzantine failures [20]. These failures, encompassing everything from intermittent data corruption to sophisticated, coordinated attacks, pose a significant threat to the convergence, accuracy, and overall stability of distributed learning algorithms. Robust distributed learning [21] aims to maintain strong learning guarantees in the presence of such failures.
In the Byzantine failure model, a misbehaving worker can send arbitrarily corrupted model updates to the server, while in data poisoning attacks, its influence is limited to manipulating local training data. Although prior work formally shows that both threat models cause similar degradation in empirical risk [15], experiments consistently find that the generalization error—i.e., expected risk on unseen data—is significantly worse under Byzantine failures [4, 26]. It remains unclear from existing literature whether this gap reflects a fundamental difference between the two threat models or the suboptimality of practical attacks. We address this open question with the first theoretical analysis of the generalization gap between Byzantine and data poisoning attacks.
To this end, we use algorithmic stability [10, 13] as our analytical tool for quantifying resilience to both forms of attacks. Algorithmic stability, which measures an algorithm’s sensitivity to changes in a single training example, offers key advantages over alternatives techniques for characterizing generalization. Unlike uniform convergence bounds [42, 7], which typically rely on i.i.d. assumptions and restricted hypothesis spaces, or PAC-Bayes [32, 6] and mutual information bounds [34], which provide less direct insight into algorithm-specific behavior, algorithmic stability takes an algorithm-centric view of generalization. This makes it particularly well-suited for robust distributed learning, where it enables a focused analysis of how specific attacks and robust aggregation rules influence the generalization performance of the overall learning procedure.
@@
This paper provides, to our knowledge, the first theoretically grounded explanation for the gap in generalization performance between two threat models in robust distributed learning: Byzantine and data poisoning attacks. Comparisons with related work are deferred to Section 5. Our main contributions—beyond formalizing a unified framework (Section 2)—are:
@@
    We derive general upper bounds on uniform stability for Byzantine attacks under convex and strongly convex objectives (Section 3.1). Specifically, for a convex loss function, our analysis shows that the uniform algorithmic stability of robust distributed $\mathrm{GD}$ and $\mathrm{SGD}$ incorporating $\mathrm{SMEA}$ [5]—an optimal high-dimensional aggregation rule— under Byzantine attacks is in $\mathcal{O}\big{(}\sqrt{\frac{f}{n-2f}}\big{)}$, $f$ being the number of misbehaving workers out of $n$.

### v2

# Byzantine Failures Harm the Generalization of Robust Distributed Learning Algorithms More Than Data Poisoning

@@ Abstract
Robust distributed learning algorithms aim to maintain reliable performance despite the presence of misbehaving workers. Such misbehaviors are commonly modeled as Byzantine failures, allowing arbitrarily corrupted communication, or as data poisoning, a weaker form of corruption restricted to local training data. While prior work shows similar optimization guarantees for both models, an important question remains: How do these threat models impact generalization? Empirical evidence suggests a gap, yet it remains unclear whether it is unavoidable or merely an artifact of suboptimal attacks. We show, for the first time, a fundamental gap in generalization guarantees between the two threat models: Byzantine failures yield strictly worse rates than those achievable under data poisoning. Our findings leverage a tight algorithmic stability analysis of robust distributed learning. Specifically, we prove that: (i) under data poisoning, the uniform algorithmic stability of an algorithm with optimal optimization guarantees degrades by an additive factor of $\varTheta(\frac{f}{n-f})$, with $f$ out of $n$ workers misbehaving; whereas (ii) under Byzantine failures, the degradation is in $\Omega\big(\sqrt{\frac{f}{n-2f}}\big)$.
@@ 1 INTRODUCTION
With the proliferation of large-scale data and pervasive connectivity, distributed learning has evolved into a fundamental paradigm in modern AI (Kairouz et al., 2021). From federated smartphones to geo-distributed datacenters, algorithms now orchestrate millions of workers toward shared objectives. Yet, unlike the idealized setting of perfectly cooperative workers, real-world systems face a range of adversarial misbehaviors, broadly categorized under the umbrella of Byzantine failures (Guerraoui et al., 2024). Encompassing everything from intermittent data corruption to sophisticated, coordinated attacks, these failures threaten the reliability of distributed learning algorithms.
Robust distributed learning aims to maintain strong learning guarantees despite the presence of such misbehaviors. The prevailing approach is to model a fraction of workers as corrupted, either Byzantine—able to send arbitrarily corrupted updates—or poisoned, where corruption is limited to local training data. While Farhadkhani et al. (2024b) showed that both threat models yield similar empirical risk guarantees, experiments consistently observe significantly worse generalization—i.e., higher risk on unseen data—under Byzantine failures (Allouah et al., 2023a; Karimireddy et al., 2022). Whether this gap reflects an inherent difference in adversarial power or only the suboptimality of practical attacks remains unknown. We address this open problem with the first theoretical analysis of the generalization gap between Byzantine failures and data poisoning.
To this end, we use algorithmic stability (Bousquet and Elisseeff, 2002) to quantify generalization resilience against both forms of attacks. In particular, uniform stability—which measures an algorithm’s sensitivity to the replacement of a single training example—offers a worst-case, distribution-free, and algorithm-centric perspective on generalization. This stands in contrast to approaches based on uniform convergence (Vapnik, 1998), PAC-Bayes (McAllester, 1999), or mutual information (Russo and Zou, 2019), which provide less direct insight into algorithm-specific behavior. This makes stability well-suited for addressing the open question under consideration, as it enables a focused analysis of the potential harm posed by different threat models on the generalization performance of robust learning algorithms.
@@
[…]
2 PRELIMINARIES
Problem setting.
We consider a distributed setup with a central server and $n$ workers, where up to $f$ (with $f<\frac{n}{2}$) may be subject to either Byzantine failures or data poisoning, as defined below. We refer to such workers as misbehaving, unless we need to distinguish between the two attack types. Their identities are unknown to the server. While the actual number of misbehaving workers may be less than $f$, we assume the worst-case scenario where exactly $f$ workers are misbehaving. The remaining honest workers form the set $\mathcal{H}\subseteq[n]\vcentcolon=\{1,\ldots,n\}$, with $|\mathcal{H}|=n-f$. Each honest worker $i\in\mathcal{H}$ holds a local dataset $\mathcal{D}_{i}=\{z^{(i,1)},\ldots,z^{(i,m)}\}$ composed of $m$ i.i.d. data points (or samples) from an input space $\mathcal{Z}$ drawn from a distribution $p_{i}$. We denote $\mathcal{S}=\cup_{i\in\mathcal{H}}\mathcal{D}_{i}$. Given a parameter vector $\theta\in\Theta\subset\mathbb{R}^{d}$ representing the model, a data point $z\in\mathcal{Z}$ incurs a differentiable loss defined by a real-valued function $\ell(\theta;z)$. The goal is to minimize the population risk over the honest workers:
As we only have access to a finite number of samples from each distribution, the above objective can only be solved approximately by minimizing the empirical risk over the honest workers, with the added challenge that the honest subgroup is unknown:
$\widehat{R}_{\mathcal{H}}(\theta)=\frac{1}{|\mathcal{H}|}\!\sum_{i\in\mathcal{H}}\widehat{R}_{i}(\theta)=\frac{1}{|\mathcal{H}|}\!\sum_{i\in\mathcal{H}}\frac{1}{m}\!\sum_{z\in\mathcal{D}_{i}}\ell(\theta,z).$
The expected excess population risk of a distributed learning algorithm’s output $\mathcal{A(\mathcal{S})}$ can be decomposed into the generalization and optimization errors (see, e.g. Hardt et al. (2016)) as follows:
$\mathbb{E}\big[R_{\mathcal{H}}(\mathcal{A(\mathcal{S})})-\inf\limits_{\theta\in\Theta}R_{\mathcal{H}}(\theta)\big]\leq\mathbb{E}\big[R_{\mathcal{H}}(\mathcal{A(\mathcal{S})})\\ -\widehat{R}_{\mathcal{H}}(\mathcal{A(\mathcal{S})})\big]+\mathbb{E}\big[\widehat{R}_{\mathcal{H}}(\mathcal{A(\mathcal{S})})-\inf\limits_{\theta\in\Theta}\widehat{R}_{\mathcal{H}}(\theta)\big]$ (1)
Prior work on robust distributed learning has focused primarily on the optimization error, i.e., the second term on the right-hand side. In contrast, our goal is to study the generalization error, corresponding to the first term. Formally, we define the notion of robustness as follows, with $\|\cdot\|_{2}$ the Euclidean norm.
Definition 2.1.
An algorithm is said to be $(f,\rho,$ statistical)-resilient if it outputs a parameter $\hat{\theta}$ s. t.
$\mathbb{E}_{\mathcal{A}}\big[R_{\mathcal{H}}(\hat{\theta})\!-\!\inf_{\theta\in\Theta}R_{\mathcal{H}}(\theta)\big]\!\leq\!\rho.$
We say $(f,\rho,\text{empirical})$-resilient when we only consider empirical risk.
Note that the (standard) notion of $(f,\rho,\text{empirical})$-resilience—see, e.g., Liu et al. (2021)—is impossible in general for any $\rho$ when $f\geq n/2$, justifying our assumption $f<n/2$.
Threat models. We aim to analyze this resilience property under two standard threat models widely studied in the literature (Farhadkhani et al., 2024a; Shejwalkar and Houmansadr, 2021).
Byzantine failures. In this threat model, misbehaving workers can act arbitrarily: they may deviate from the prescribed algorithm, collude, and send adversarial updates to the server while having access to all information exchanged between honest workers and the server. This model follows the classical notion of Byzantine failures in distributed systems introduced in the seminal work of Lamport et al. (1982).
Data poisoning. In this restricted threat model, misbehaving workers follow the prescribed algorithm but may corrupt their local training data before execution. Specifically, for each misbehaving worker $i\notin\mathcal{H}$, a local dataset $\mathcal{D}_{i}\in\mathcal{Z}^{m}$ is adversarially constructed with full knowledge of the honest workers’ data distributions $\{p_{i},~i\in\mathcal{H}\}$.
Background on robust distributed optimization.
Robust distributed optimization algorithms aim to achieve $(f,\rho,\text{empirical})$-resilience. They are typically adaptations of standard first-order iterative optimization algorithms like gradient descent ($\mathrm{GD}$) or stochastic gradient descent ($\mathrm{SGD}$). These algorithms are made resilient by replacing the server-side averaging operator by a robust aggregation rule $F:(\mathbb{R}^{d})^{n}\to\mathbb{R}^{d}$ (Guerraoui et al., 2024). Formally, given a learning rate $\gamma$, the parameter $\theta_{t}$ at iteration $t\in\{0,\ldots,T-1\}$ is updated as follows:
where $g^{(i)}_{t}$ denotes the update sent by worker $i$ at iteration $t$. For an honest worker $i\in\mathcal{H}$, we have $g^{(i)}_{t}=\nabla\widehat{R}_{i}(\theta_{t})$ under $\mathrm{GD}$, or $g^{(i)}_{t}=\nabla\ell(\theta_{t};z^{(i)}_{t})$ under $\mathrm{SGD}$, with $z^{(i)}_{t}$ uniformly sampled from the local dataset $\mathcal{D}_{i}$. For a misbehaving worker $i\notin\mathcal{H}$, $g^{(i)}_{t}$ is either an arbitrary vector in $\mathbb{R}^{d}$ (Byzantine failures) or a gradient computed on corrupted data (data poisoning).
Recent advances in robust distributed optimization have identified key properties sufficient for a robust aggregation rule to ensure $(f,\rho,\text{empirical})$-resilience (Allouah et al., 2023a, b). The following definition encompasses a wide range of aggregation rules and has been shown to yield tight resilience guarantees across a variety of practical distributed learning settings.
Definition 2.2.
Let $n\!\geq\!1$, $0\!\leq f\!<\!n/2$ and $\kappa\!\geq\!0$. An aggregation rule $F$ is $(f,\kappa)$-robust if for any $g_{1},\ldots,g_{n}\in\mathbb{R}^{d}$, any set $S\!\subset\![n]$ of size nfn-f, with $\overline{g}_{S}=\frac{1}{|S|}\sum_{i\in S}g_{i}$ and $\Sigma_{S}=\frac{1}{|S|}\sum_{i\in S}(g_{i}-\overline{g}_{S}){(g_{i}-\overline{g}_{S})}^{\intercal}$,
$\|F(g_{1},\ldots,g_{n})-\overline{g}_{S}\|_{2}^{2}\leq\kappa\|\Sigma_{S}\|_{sp}.$
Here, $\|\cdot\|_{sp}$ denotes the spectral norm (i.e., the largest eigenvalue), $f$ and $\kappa$ are referred to as the robustness parameter and robustness coefficient of $F$, respectively.
Any aggregation rule $F$ that is $(f,\mathcal{O}(f/n))$-robust achieves optimal $(f,\rho,\text{empirical})$-resilience. In particular, iterative methods employing such a rule attain an optimization error that matches the information-theoretic lower bound for first-order optimization in the presence of $f$ Byzantine workers. One such optimally robust rule is SMEA (Allouah et al., 2023b).
Definition 2.3.
Given $f,n\in\mathbb{N}$, $f<n/2$ and vectors $g_{1},\ldots,g_{n}\in\mathbb{R}^{d}$, the smallest maximum eigenvalue averaging aggregation rule (SMEA) outputs the average of the nfn-f gradients that are most directionally consistent, discarding outliers based on covariance:
$\overline{g}_{S^{*}}=\frac{1}{|S^{*}|}\sum_{i\in S^{*}_{t}}g_{i},\quad\text{with}\quad S^{*}\in\underset{\begin{subarray}{c}{S\subseteq\{1,\ldots,n\},}\\ {|S|=n-f}\end{subarray}}{\arg\!\min}\|\Sigma_{S}\|_{sp}$
However, the above results focus on the optimization error, overlooking the impact of robust aggregation rules on the generalization error—the first term in the right-hand side of eq. 1. To address this gap, we provide a rigorous analysis of the generalization guarantees of robust distributed learning, focusing on the $\mathrm{SMEA}$ aggregation rule as it yields optimal optimization guarantees.
Stability under misbehaving workers.
Our study leverages the framework of algorithmic stability to understand the generalization error of robust distributed learning algorithms. We do so via uniform algorithmic stability, which measures how sensitive a learning algorithm is to changes in its training data. This approach was introduced by Vapnik and Chervonenkis (1974); Devroye and Wagner (1979); Rogers and Wagner (1978) and later popularized by Bousquet and Elisseeff (2002); Shalev-Shwartz et al. (2010); Hardt et al. (2016). We revisit this framework to study generalization in robust distributed learning, where the change is only in the honest workers’ data.
Definition 2.4.
Consider $n$ workers, with $f$ misbehaving and nfn-f honest, each holding $m$ local samples. A distributed algorithm $\mathcal{A}$ is $\varepsilon$-uniformly stable if, for all honest datasets $\mathcal{S},\mathcal{S}^{\prime}\in\mathcal{Z}^{(n-f)m}$ that are neighboring—i.e., differing in at most one sample—we have
$\sup_{z\in\mathcal{Z}}\mathbb{E}[\ell(\mathcal{A}(\mathcal{S});z)-\ell(\mathcal{A}(\mathcal{S}^{\prime});z)]\leq\varepsilon,$
where the expectation is over the randomness of $\mathcal{A}$.
This property leads to the following bound on the generalization error, proved in Appendix A.
Proposition 2.1.
If $\mathcal{A}$ is $\varepsilon$-uniformly stable, then
The link between uniform stability and generalization can be further strengthened to yield high-probability generalization bounds, as discussed in Appendix G. We can now formalize a relationship between empirical resilience and statistical resilience through stability: any algorithm that is $(f,\rho,\text{empirical})$-resilient and $\varepsilon$-uniformly stable is $(f,\rho+\varepsilon,\text{statistical})$-resilient. Here, $\varepsilon$ quantifies the effect of misbehaving workers on generalization, and is the central focus of this paper.
3 (IN)STABILITY OF ROBUST DISTRIBUTED LEARNING
In this section, we first derive upper bounds on the stability of robust distributed $(\mathrm{S})\mathrm{GD}$ under general $(f,\kappa)$-robust aggregation rules. We next analyze $\mathrm{SMEA}$ to derive lower bounds under Byzantine failures, as well as tight upper and lower bounds for the special case of data poisoning. Our results show a fundamental gap in stability between the two threat models. Definitions of function regularity are given in Appendix B.
3.1 The Case of Byzantine Failures
Upper bounds. We start by deriving uniform stability upper bounds for convex optimization. Proofs are deferred to Section C.2.
Theorem 3.1.
Consider the setting described in Section 2 under Byzantine failures. Let $\mathcal{A}\in\{\mathrm{GD},\mathrm{SGD}\}$ with a $(f,\kappa)$-robust aggregation rule $F$. Suppose $\forall z\in\mathcal{Z}$, $\ell(\cdot;z)$ $C$-Lipschitz and $L$-smooth, and $\mathcal{A}$ is run for $T\in\mathbb{N}^{*}$ iterations with $\gamma\leq\frac{1}{L}$.
@@ 3.1 The Case of Byzantine Failures
    If $\ell(\cdot;z)$ is convex $\forall z\in\mathcal{Z}$, then the uniform stability of $\mathcal{A}$ is upper bounded by
    $2\gamma C^{2}T\left(\sqrt{\kappa}+\frac{1}{(n-f)m}\right).$ (2)